# Pauli Hamiltonians with an Aharonov-Bohm flux

### v1

@@ 1. Introduction
$\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\,\qquad\sigma_{2}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}.$ (1.2)
@@
It is known that the singularity at the origin of the field affects the self-adjointness of the Hamiltonian operator and, accordingly, different self-adjoint extensions of the formal expression exist. The analogous subject for the Schrödinger operator, i.e. for a spinless charged particle, has been thoroughly analysed in the literature (we refer to [AT98, CF20, CF23, DS98, F23, PR11] for further details; see also [DF23, DG21] for results about the related radial operators) and our starting point is a similar classification of all self-adjoint realizations of the Pauli operator (1.1) (§ 2.1 and § 2.2). Such a question has only been partially addressed in [GS04, Pe05, Pe06], also for more than one solenoid, yet focusing mainly on the Aharonov-Casher phenomenon about the existence and number of zero-energy modes. In this connection we also mention [EV02], where a distinguished self-adjoint extension if characterized for Pauli operators with measure-valued magnetic fields.
@@
Acknowledgments. The present research has been supported by MUR grant “Dipartimento di Eccellenza” 2023-2027 of Dipartimento di Matematica, Politecnico di Milano. MC acknowledges the support of PNRR Italia Domani and Next Generation Eu through the ICSC National Research Centre for High Performance Computing, Big Data and Quantum Computing. WB is supported by the PRIN 2022 Project “Nonlinear effective equations in presence of singularities” and by the GNAMPA 2023 Project “Analisi spettrale per equazioni di Dirac e applicazioni”.
@@ 2. Main Results
Let us stress that $U$ is indeed a unitary map, yet not a gauge transformation since the magnetic fluxes associated to $H_{\mathrm{P}}$ and $UH_{\mathrm{P}}U^{-1}$ are different. The operator $H_{\mathrm{P}}$ is understood as a closable symmetric operator on the dense domain $C^{\infty}_{c}(\mathbb{R}^{2}\setminus\{\mathbf{0}\};\mathbb{C}^{2})=C^{\infty}_{c}(\mathbb{R}^{2}\setminus\{\mathbf{0}\})\otimes\mathbb{C}^{2}$. However, it is also easy to verify that $H_{\mathrm{P}}$ is not essentially self-adjoint on such a domain, due to the singularity of $\mathbf{A}$ at the origin. At the algebraic level, the operator (1.1) can be written as
@@
where $\Pi_{\pm}:=(p_{1}\pm ip_{2})$ and $p_{j}$, $j=1,2$, are the components of the vector $\mathbf{p}:=-i\nabla+\mathbf{A}$. It is readily seen that $\Pi_{\pm}$ are one the formal adjoint of the other. Passing to polar coordinates and noting that $\mathbf{A}(r,\theta)=\tfrac{\alpha}{r}\,(-\sin\theta,\cos\theta)$, $\partial_{x}=\cos\theta\,\partial_{r}-\tfrac{\sin\theta}{r}\,\partial_{\theta}$ and $\partial_{y}=\sin\theta\,\partial_{r}+\tfrac{\cos\theta}{r}\,\partial_{\theta}$, by a simple computation we get
$\Pi_{+}=e^{i\theta}\left(-i\partial_{r}+\tfrac{1}{r}\,\partial_{\theta}+\tfrac{i\alpha}{r}\right),\qquad\Pi_{-}=e^{-i\theta}\left(-i\partial_{r}-\tfrac{1}{r}\,\partial_{\theta}-\tfrac{i\alpha}{r}\right).$
Then, using the basic relation $\big{[}\partial_{\theta},e^{\pm i\theta}\big{]}=\pm i\,e^{i\theta}$, we find
$\Pi_{-}\Pi_{+}$ $=\left(-i\partial_{r}-\tfrac{1}{r}\,\partial_{\theta}-\tfrac{i(\alpha+1)}{r}\right)\left(-i\partial_{r}+\tfrac{1}{r}\,\partial_{\theta}+\tfrac{i\alpha}{r}\right)=-\,\tfrac{1}{r}\,\partial_{r}\big{(}r\,\partial_{r}\,\cdot\big{)}+\tfrac{1}{r^{2}}\,(-i\,\partial_{\theta}+\alpha)^{2}\,,$
$\Pi_{+}\Pi_{-}$ $=\left(-i\partial_{r}+\tfrac{1}{r}\,\partial_{\theta}+\tfrac{i(\alpha-1)}{r}\right)\left(-i\partial_{r}-\tfrac{1}{r}\,\partial_{\theta}-\tfrac{i\alpha}{r}\right)=-\,\tfrac{1}{r}\,\partial_{r}\big{(}r\,\partial_{r}\,\cdot\big{)}+\tfrac{1}{r^{2}}\,(-i\,\partial_{\theta}+\alpha)^{2}\,.$
@@ 2.1. Friedrichs extension
$\left\langle f\right\rangle(r):=\frac{1}{2\pi}\int_{0}^{2\pi}\!\!\!\mathrm{d}\vartheta\>f(r,\theta)\,.$
@@ 2.2. Self-adjoint extensions
$g_{\lambda}^{(\ell)}(r,\theta)=\;\lambda^{|\ell+\alpha|}\,K_{|\ell+\alpha|}(\lambda r)\,\tfrac{e^{i\ell\theta}}{\sqrt{2\pi}},\qquad\ell\in\{0,-1\}\,.$ (2.8)
@@
$\psi_{s}(r,\theta)=\sum_{\ell\in\mathbb{Z}}\,\psi_{s}^{(\ell)}(r)\,\tfrac{e^{i\ell\theta}}{\sqrt{2\pi}}\,.$
@@
The Friedrichs Hamiltonian $H_{\mathrm{P}}^{(\mathrm{F})}$ is formally recovered taking $\beta=\mbox{``}\infty\mbox{''}\mathds{1}$, i.e., setting all the charges $\mathbf{q}$ equal to zero. Another notable extension is the Krein’s one, i.e., that with the largest possible domain. Here, it is easier to identify it as the unique extension besides the Friedrichs’ one which is scale invariant, namely, for given $\lambda>0$, it coincides with $H_{\mathrm{P}}^{(\beta_{\lambda})}$, where $\beta_{\lambda}=-L(\lambda)$.
@@
$H_{\mathrm{P}}^{(U)}\bm{\psi}=H_{\mathrm{P}}\bm{f}+i\,\mathcal{G}_{+}\mathbf{c}-i\,\mathcal{G}_{+}U\mathbf{c}\,,$ (2.16)
@@
$g_{\pm}^{(\ell)}(r,\theta)=e^{\mp i\tfrac{\pi}{4}\,|\ell+\alpha|}\,\sqrt{\tfrac{4}{\pi}\cos\big{(}\tfrac{\pi}{2}\,|\alpha+\ell|\big{)}}\,\textstyle{e^{i\ell\theta}\over\sqrt{2\pi}}\,K_{|\ell+\alpha|}\big{(}e^{\mp i\pi/4}\,r\big{)}\,,\qquad\ell\in\{0,-1\}\,.$ (2.19)
@@ 2.3. Spectral and scattering properties
$R_{\mathrm{S}}^{(\mathrm{F})}(z;\mathbf{x},\mathbf{x}^{\prime})=\sum_{\ell\in\mathbb{Z}}\,I_{|\ell+\alpha|}\big{(}-i\sqrt{z}\,(r\wedge r^{\prime})\big{)}\,K_{|\ell+\alpha|}\big{(}-i\sqrt{z}\,(r\land r^{\prime})\big{)}\,\tfrac{e^{i\ell(\theta-\theta^{\prime})}}{2\pi}\,.$ (2.21)
@@
There is a one-to-one correspondence between the families $\big{\{}H_{\mathrm{P}}^{(\Theta)}\big{\}}_{\Theta\in M_{4,\mathrm{Herm}}(\mathbb{C})}$ and $\big{\{}H_{\mathrm{P}}^{(\beta)}\big{\}}_{\beta\in M_{4,\mathrm{Herm}}(\mathbb{C})}$, given for any fixed $\lambda$ large enough by
$\Theta=\Theta(\beta)=L(\lambda)+\beta\,,$ (2.28)
where $L(\lambda)$ is defined in (2.12).
@@
Also in this case it appears that the Friedrichs Hamiltonian $H_{\mathrm{P}}^{(\mathrm{F})}$ is formally recovered for $\Theta=\mbox{``}\infty\mbox{''}\mathds{1}$, while the Krein’s one is simply identified by $\Theta=0$.
@@
where $\widehat{U}$ is a unitary operator on the defect space $\mathrm{span}\{\mathcal{G}_{+}\mathbf{c}\}$, $\mathbf{c}\in\mathbb{C}^{4}$, acting as $\widehat{U}\,\mathcal{G}_{+}\mathbf{c}:=\mathcal{G}_{-}\,U\mathbf{c}$, and $\widehat{U}_{\ast}:=\big{(}H_{\mathrm{P}}^{(\mathrm{F})}\!-i\big{)}R_{\mathrm{P}}^{(\mathrm{F})}(-i)$ is the restriction of the Cayley transform of $H_{\mathrm{P}}^{(\mathrm{F})}$ to the same subspace. Of course the family $H_{\mathrm{P}}^{(U)}$, $U\in\mathrm{M}_{4,\mathrm{Unit}}(\mathbb{C})$, comprises the Friedrichs and Krein realizations, which are respectively recovered for
$U^{(\mathrm{F})}=-\mathds{1}\,,\qquad U^{(\mathrm{K})}=\mathds{1}\,.$
@@
with domain $H^{2}(\mathbb{R}^{2};\mathbb{C}^{2})$. Notice that $\sigma(-\Delta_{\mathrm{P}})=\sigma_{\mathrm{ac}}(-\Delta_{P})=\mathbb{R}^{+}$, so the projector onto the subspace of absolute continuity of $-\Delta_{P}$ satisfies $P_{\mathrm{ac}}(-\Delta_{P})=\mathds{1}$. More in general, for any $\Theta\in\mathrm{M}_{4,\mathrm{Herm}}(\mathbb{C})$, we define the wave operators
@@
Notice that $\mathrm{S}^{(\Theta)}$ is a unitary operator on $\mbox{ran}\,P_{\mathrm{ac}}\big{(}H_{\mathrm{P}}^{(\Theta)}\big{)}$ as soon as the wave operator $\Omega^{(\Theta)}_{\pm}$ are complete. Finally, we recall that a zero-energy resonance of $H_{\mathrm{P}}^{(\Theta)}$ is a distributional solution of the equation $H_{\mathrm{P}}\bm{\psi}=0$ in $L^{2}_{\mathrm{loc}}(\mathbb{R}^{2})\setminus L^{2}(\mathbb{R}^{2})$, which fulfils the boundary condition at $\mathbf{x}=\mathbf{0}$ encoded in $\mathscr{D}\big{(}H_{\mathrm{P}}^{(\Theta)}\big{)}$ and remains bounded at infinity.
@@
The wave operators $\Omega^{(\mathrm{F})}_{\pm}$ exists and are asymptotically complete. Moreover, the scattering operator $\mathrm{S}^{(\mathrm{F})}$ exists and is unitary on $L^{2}(\mathbb{R}^{2};\mathbb{C}^{2})$.
@@
$\bm{\varphi}_{(s,\mathbf{k})}(\mathbf{x}):=\tfrac{1}{2\pi}\,e^{i\mathbf{k}\cdot\mathbf{x}}\begin{pmatrix}\delta_{s,\uparrow}\\ \delta_{s,\downarrow}\end{pmatrix}=\tfrac{1}{2\pi}\textstyle\sum_{\ell}e^{i\ell(\theta-\omega)+i\frac{\pi}{2}\,|\ell|}\,J_{|\ell|}(k\,r)\,\begin{pmatrix}\delta_{s,\uparrow}\\ \delta_{s,\downarrow}\end{pmatrix},$ (2.31)
where $\mathbf{x}=(r,\theta)\in\mathbb{R}^{+}\!\times\!\mathbb{S}^{1}$, $\mathbf{k}=(k,\omega)\in\mathbb{R}^{+}\!\times\!\mathbb{S}^{1}$ and we used [GR07, Eq. 8.511.4] and [OLBC10, Eq. 10.4.1]. Notice that, in the sense of distributions,
@@
$\bm{\varphi}^{(\mathrm{F},\pm)}_{(s,\mathbf{k})}(\mathbf{x})=\tfrac{1}{2\pi}\sum_{\ell\in\mathbb{Z}}e^{i\ell(\theta-\omega_{\pm})\pm i\frac{\pi}{2}\,|\ell+\alpha|}\,J_{|\ell+\alpha|}(k\,r)\,\begin{pmatrix}\delta_{s,\uparrow}\\ \delta_{s,\downarrow}\end{pmatrix},$ (2.37)
@@
We do not discuss here the presence of eigenvalues embedded in the continuous spectrum. These would be exactly the exceptional points forming the set $\mathrm{e}_{+}^{(\Theta)}$ identified next in Proposition 3.3. Such a question has been investigated for regular magnetic fields in [HH23, AHH20]
@@
$f^{(\Theta,\pm)}_{(s,s^{\prime}),\mathbf{k}}:=\tfrac{e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell\in\mathbb{Z}}\left(e^{\pm i\pi|\ell+\alpha|}-e^{\pm i\pi|\ell|}\right)e^{i\ell(\theta-\omega_{\pm})}\,\delta_{s\,s^{\prime}}\\ +\tfrac{i\pi\,e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell,\ell^{\prime}\in\{0,-1\}}\!\Big{\{}\big{[}\Lambda_{\pm}(k^{2})+\Theta\big{]}^{-1}\Big{\}}_{s^{\prime}\!,s}^{\ell^{\prime}\!,\ell}\,e^{i(\ell^{\prime}\theta-\ell\omega_{\pm})}\;(\pm ik)^{|\ell+\alpha|+|\ell^{\prime}+\alpha|}\,.$ (2.46)
@@
Proposition 2.19 (Zero-energy resonances of $H_{\mathrm{P}}^{(F)}$ and $H_{\mathrm{P}}^{(\Theta)}$).
@@
$\bm{\psi}_{0}=\begin{pmatrix}\psi_{\uparrow}\\ \psi_{\downarrow}\end{pmatrix},\qquad\psi_{s}(r,\theta)=\sum_{\ell\in\{0,-1\}}q_{s}^{(\ell)}\,\tfrac{2^{|\ell+\alpha|-1}\Gamma(|\ell+\alpha|)}{r^{|\ell+\alpha|}}\,\tfrac{e^{i\ell\theta}}{\sqrt{2\pi}},$ (2.47)

### v3

@@ 1. Introduction
$\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\,,\qquad\sigma_{2}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}.$ (1.2)
@@
It is known that the singularity at the origin of the field affects the self-adjointness of the Hamiltonian operator and, accordingly, different self-adjoint extensions of the formal expression exist. The analogous subject for the Schrödinger operator, i.e. for a spinless charged particle, has been thoroughly analysed in the literature (we refer to [AT98, CF20, CF23, CF24, DS98, F23, PR11] for further details; see also [DF23, DG21] and [GHT23, GPS22] for results about the related radial operators) and our starting point is a similar classification of all self-adjoint realizations of the Pauli operator (1.1) (§ 2.1 and § 2.2). Such a question has only been partially addressed in [GS04, Pe05, Pe06], also for more than one solenoid, yet focusing mainly on the Aharonov-Casher phenomenon about the existence and number of zero-energy modes. In this connection we also mention [EV02], where a distinguished self-adjoint extension is characterized for Pauli operators with measure-valued magnetic fields.
@@
Acknowledgments. The present research has been supported by MUR grant “Dipartimento di Eccellenza” 2023-2027 of Dipartimento di Matematica, Politecnico di Milano. MC acknowledges the support of PNRR Italia Domani and Next Generation Eu through the ICSC National Research Centre for High Performance Computing, Big Data and Quantum Computing. WB acknowledges the support of the MUR - PRIN 2022 project “Nonlinear dispersive equations in presence of singularities” (Prot. N. 20225ATSTP).
@@ 2. Main Results
Let us stress that $U$ is indeed a unitary map, yet not a gauge transformation since the magnetic fluxes associated to $H_{\mathrm{P}}$ and $UH_{\mathrm{P}}U^{-1}$ are different.
The operator $H_{\mathrm{P}}$ is understood as a closable symmetric operator on the dense domain $C^{\infty}_{c}(\mathbb{R}^{2}\setminus\{\mathbf{0}\};\mathbb{C}^{2})=C^{\infty}_{c}(\mathbb{R}^{2}\setminus\{\mathbf{0}\})\otimes\mathbb{C}^{2}$. However, it is also easy to verify that $H_{\mathrm{P}}$ is not essentially self-adjoint on such a domain, due to the singularity of $\mathbf{A}$ at the origin. At the algebraic level, the operator (1.1) can be written as
@@
where $\Pi_{\pm}:=(p_{1}\pm ip_{2})$ and $p_{j}$, $j=1,2$, are the components of the vector $\mathbf{p}:=-i\nabla+\mathbf{A}$. It is readily seen that $\Pi_{\pm}$ are one the formal adjoint of the other. Passing to polar coordinates and noting that $\mathbf{A}(r,\vartheta)=\tfrac{\alpha}{r}\,(-\sin\vartheta,\cos\vartheta)$, $\partial_{x}=\cos\vartheta\,\partial_{r}-\tfrac{\sin\vartheta}{r}\,\partial_{\vartheta}$ and $\partial_{y}=\sin\vartheta\,\partial_{r}+\tfrac{\cos\vartheta}{r}\,\partial_{\vartheta}$, by a simple computation we get
$\Pi_{+}=e^{i\vartheta}\left(-i\partial_{r}+\tfrac{1}{r}\,\partial_{\vartheta}+\tfrac{i\alpha}{r}\right),\qquad\Pi_{-}=e^{-i\vartheta}\left(-i\partial_{r}-\tfrac{1}{r}\,\partial_{\vartheta}-\tfrac{i\alpha}{r}\right).$
Then, using the basic relation $\big{[}\partial_{\vartheta},e^{\pm i\vartheta}\big{]}=\pm i\,e^{i\vartheta}$, we find
$\Pi_{-}\Pi_{+}$ $=\left(-i\partial_{r}-\tfrac{1}{r}\,\partial_{\vartheta}-\tfrac{i(\alpha+1)}{r}\right)\left(-i\partial_{r}+\tfrac{1}{r}\,\partial_{\vartheta}+\tfrac{i\alpha}{r}\right)=-\,\tfrac{1}{r}\,\partial_{r}\big{(}r\,\partial_{r}\,\cdot\big{)}+\tfrac{1}{r^{2}}\,(-i\,\partial_{\vartheta}+\alpha)^{2}\,,$
$\Pi_{+}\Pi_{-}$ $=\left(-i\partial_{r}+\tfrac{1}{r}\,\partial_{\vartheta}+\tfrac{i(\alpha-1)}{r}\right)\left(-i\partial_{r}-\tfrac{1}{r}\,\partial_{\vartheta}-\tfrac{i\alpha}{r}\right)=-\,\tfrac{1}{r}\,\partial_{r}\big{(}r\,\partial_{r}\,\cdot\big{)}+\tfrac{1}{r^{2}}\,(-i\,\partial_{\vartheta}+\alpha)^{2}\,.$
@@ 2.1. Friedrichs extension
$\left\langle f\right\rangle(r):=\frac{1}{2\pi}\int_{0}^{2\pi}\!\!\!\mathrm{d}\vartheta\>f(r,\vartheta)\,.$
@@ 2.2. Self-adjoint extensions
$g_{\lambda}^{(\ell)}(r,\vartheta)=\;\lambda^{|\ell+\alpha|}\,K_{|\ell+\alpha|}(\lambda r)\,\tfrac{e^{i\ell\vartheta}}{\sqrt{2\pi}},\qquad\ell\in\{0,-1\}\,.$ (2.8)
@@
$\psi_{s}(r,\vartheta)=\sum_{\ell\in\mathbb{Z}}\,\psi_{s}^{(\ell)}(r)\,\tfrac{e^{i\ell\vartheta}}{\sqrt{2\pi}}\,.$
@@
Remark 2.4 (Local singularities).
For a generic $\beta\in M_{4,\,\mathrm{Herm}}(\mathbb{C})$, considering that the Friedrichs Hamiltonian $H_{\mathrm{P}}^{(\mathrm{F})}$ is the direct sum of two copies of the Friedrichs AB Schrödinger operator (see Remark 2.2), working in polar coordinates, using results from [DR17] and exploiting the asymptotic expansion of $\bm{G}_{\lambda,s}^{(\ell)}$ (see (2.9) and (3.2), below), for any $\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{P}}^{(\beta)}\big{)}$ one infers that, as $r\to 0^{+}$,
$\psi_{s}(r,\theta)=\textstyle{\sum_{\ell\,\in\,\{0,-1\}}}\left[\tfrac{\Gamma(|\ell+\alpha|)}{2^{1-|\ell+\alpha|}}\,\tfrac{q_{s}^{(\ell)}}{r^{|\ell+\alpha|}}+\left(\phi_{s}^{(\ell)}+q_{s}^{(\ell)}\,\tfrac{\Gamma(\!-|\ell+\alpha|)}{2^{1+|\ell+\alpha|}}\,\lambda^{2|\ell+\alpha|}\right)r^{|\ell+\alpha|}\right]\tfrac{e^{i\ell\theta}}{\sqrt{2\pi}}+o(r)\,,$
for some $\phi_{s}^{(\ell)}\in\mathbb{C}$. Noticing that the boundary conditions in (2.13) can be equivalently expressed as
$\big{[}\big{(}L(\lambda)+\beta\big{)}\mathbf{q}\big{]}_{s}^{(\ell)}=2^{|\ell+\alpha|}\,\Gamma\big{(}|\ell+\alpha|+1\big{)}\,\phi_{s}^{(\ell)}\,,$
and recalling the definition (2.12) of $L(\lambda)$, the above asymptotics can be equivalently expressed as
$\psi_{s}(r,\theta)=\textstyle{\sum_{\ell\,\in\,\{0,-1\}}}\bigg{[}\tfrac{\Gamma(|\ell+\alpha|)}{2^{1-|\ell+\alpha|}}\,{q_{s}^{(\ell)}\over r^{|\ell+\alpha|}}+\frac{1}{2^{|\ell+\alpha|}\,\Gamma(|\ell+\alpha|+1)}\,\big{[}\beta\mathbf{q}\big{]}_{s}^{(\ell)}\,r^{|\ell+\alpha|}\bigg{]}\,\tfrac{e^{i\ell\theta}}{\sqrt{2\pi}}+o(r)\,,$ (2.15)
@@
The Friedrichs Hamiltonian $H_{\mathrm{P}}^{(\mathrm{F})}$ is formally recovered taking $\beta=\mbox{``}\infty\mbox{''}\mathds{1}$, i.e., setting all the charges $\mathbf{q}$ equal to zero. Another notable extension is the Krein’s Hamiltonian, i.e., the smallest positive one in form sense. By directed inspection of (2.11), noting that $L(\lambda)$ is positive (see (2.12)), it can be checked that the latter coincides with $H_{\mathrm{P}}^{(0)}$, namely, the extension with $\beta=0$. The same conclusion can be inferred by considering the asymptotics (2.15) and identifying the Krein’s Hamiltonian as the unique extension, besides the Friedrichs’ one, which is homogeneous of degree $-2$ under the $L^{2}$-unitary scaling
$u(\cdot)\mapsto\eta\,u(\eta\,\cdot)\qquad(\eta>0)\,.$
@@
$H_{\mathrm{P}}^{(U)}\bm{\psi}=H_{\mathrm{P}}\bm{f}+i\,\mathcal{G}_{+}\mathbf{c}-i\,\mathcal{G}_{-}U\mathbf{c}\,,$ (2.17)
@@
$g_{\pm}^{(\ell)}(r,\vartheta)=e^{\mp i\tfrac{\pi}{4}\,|\ell+\alpha|}\,\sqrt{\tfrac{4}{\pi}\cos\big{(}\tfrac{\pi}{2}\,|\alpha+\ell|\big{)}}\,\textstyle{e^{i\ell\vartheta}\over\sqrt{2\pi}}\,K_{|\ell+\alpha|}\big{(}e^{\mp i\pi/4}\,r\big{)}\,,\qquad\ell\in\{0,-1\}\,.$ (2.20)
@@
Comparing the asymptotics at the origin (2.15) for the functions in the domains $\mathscr{D}\big{(}H_{\mathrm{P}}^{(\beta)}\big{)}$ and $\mathscr{D}\big{(}H_{\mathrm{P}}^{(U)}\big{)}$ (see (2.13) and (2.16), respectively) one can check that the Friedrichs and Krein realizations, respectively, correspond to
$U^{(\mathrm{F})}=-\mathds{1}\,,\qquad U^{(\mathrm{K})}=\big{(}-e^{i\pi|\alpha+\ell|}\delta_{ss^{\prime}}\delta_{\ell\ell^{\prime}}\big{)}\,.$
@@ 2.3. Spectral and scattering properties
$R_{\mathrm{S}}^{(\mathrm{F})}(z;\mathbf{x},\mathbf{x}^{\prime})=\sum_{\ell\in\mathbb{Z}}\,I_{|\ell+\alpha|}\big{(}-i\sqrt{z}\,(r\land r^{\prime})\big{)}\,K_{|\ell+\alpha|}\big{(}-i\sqrt{z}\,(r\lor r^{\prime})\big{)}\,\tfrac{e^{i\ell(\vartheta-\vartheta^{\prime})}}{2\pi}\,.$ (2.22)
@@
There is a one-to-one correspondence between the families $\big{\{}H_{\mathrm{P}}^{(\Theta)}\big{\}}_{\Theta\in M_{4,\mathrm{Herm}}(\mathbb{C})}$ and $\big{\{}H_{\mathrm{P}}^{(\beta)}\big{\}}_{\beta\in M_{4,\mathrm{Herm}}(\mathbb{C})}$, given by
$\Theta=\Theta(\beta)=L(1)+\beta\,,$ (2.29)
[…]
@@
Also in this case it appears that the Friedrichs Hamiltonian $H_{\mathrm{P}}^{(\mathrm{F})}$ is formally recovered for $\Theta=\mbox{``}\infty\mbox{''}\mathds{1}$, while the Krein’s one is identified by $\Theta=\Theta(0)=L(1)$.
@@
where $\widehat{U}$ is a unitary operator on the defect space $\mathrm{span}\{\mathcal{G}_{+}\mathbf{c}\}$, $\mathbf{c}\in\mathbb{C}^{4}$, acting as $\widehat{U}\,\mathcal{G}_{+}\mathbf{c}:=\mathcal{G}_{-}\,U\mathbf{c}$, and $\widehat{U}_{\ast}:=\big{(}H_{\mathrm{P}}^{(\mathrm{F})}\!-i\big{)}R_{\mathrm{P}}^{(\mathrm{F})}(-i)$ is the restriction of the Cayley transform of $H_{\mathrm{P}}^{(\mathrm{F})}$ to the same subspace.
@@
with domain $H^{2}(\mathbb{R}^{2};\mathbb{C}^{2})$. Notice that $\sigma(-\Delta_{\mathrm{P}})=\sigma_{\mathrm{ac}}(-\Delta_{\mathrm{P}})=\mathbb{R}^{+}$, so the projector onto the subspace of absolute continuity of $-\Delta_{\mathrm{P}}$ satisfies $P_{\mathrm{ac}}(-\Delta_{\mathrm{P}})=\mathds{1}$. More in general, for any $\Theta\in\mathrm{M}_{4,\mathrm{Herm}}(\mathbb{C})$, we define the wave operators
@@
Notice that $\mathrm{S}^{(\Theta)}$ is a unitary operator on $\mbox{ran}\,P_{\mathrm{ac}}\big{(}H_{\mathrm{P}}^{(\Theta)}\big{)}$ as soon as the wave operator $\Omega^{(\Theta)}_{\pm}$ are complete.
Finally, we introduce the following definition of zero-energy resonances of Pauli operators. We adopt an analogous definition for the zero-energy resonances of the Dirac operator (see next Proposition 2.24 and Proposition 2.27).
Definition 2.1 (Zero-energy resonance).
A zero-energy resonance $\bm{\psi}$ of $H_{\mathrm{P}}^{(\Theta)}$ is a distributional solution of the equation $H_{\mathrm{P}}\bm{\psi}=0$ in $L^{2}_{\mathrm{loc}}(\mathbb{R}^{2})\setminus L^{2}(\mathbb{R}^{2})$, which fulfils the boundary condition at $\mathbf{x}=\mathbf{0}$ encoded in $\mathscr{D}\big{(}H_{\mathrm{P}}^{(\Theta)}\big{)}$ and remains bounded at infinity.
@@
The wave operators $\Omega^{(\mathrm{F})}_{\pm}$ exist and are asymptotically complete. Moreover, the scattering operator $\mathrm{S}^{(\mathrm{F})}$ exists and is unitary on $L^{2}(\mathbb{R}^{2};\mathbb{C}^{2})$.
@@
$\bm{\varphi}_{(s,\mathbf{k})}(\mathbf{x}):=\tfrac{1}{2\pi}\,e^{i\mathbf{k}\cdot\mathbf{x}}\begin{pmatrix}\delta_{s,\uparrow}\\ \delta_{s,\downarrow}\end{pmatrix}=\tfrac{1}{2\pi}\textstyle\sum_{\ell}e^{i\ell(\vartheta-\omega)+i\frac{\pi}{2}\,|\ell|}\,J_{|\ell|}(k\,r)\,\begin{pmatrix}\delta_{s,\uparrow}\\ \delta_{s,\downarrow}\end{pmatrix},$ (2.32)
where $\mathbf{x}=(r,\vartheta)\in\mathbb{R}^{+}\!\times\!\mathbb{S}^{1}$, $\mathbf{k}=(k,\omega)\in\mathbb{R}^{+}\!\times\!\mathbb{S}^{1}$ and we used [GR07, Eq. 8.511.4] and [OLBC10, Eq. 10.4.1]. Notice that, in the sense of distributions,
@@
$\bm{\varphi}^{(\mathrm{F},\pm)}_{(s,\mathbf{k})}(\mathbf{x})=\tfrac{1}{2\pi}\sum_{\ell\in\mathbb{Z}}e^{i\ell(\vartheta-\omega_{\pm})\pm i\frac{\pi}{2}\,|\ell+\alpha|}\,J_{|\ell+\alpha|}(k\,r)\,\begin{pmatrix}\delta_{s,\uparrow}\\ \delta_{s,\downarrow}\end{pmatrix},$ (2.38)
@@
Remark 2.18 (An extension with negative point spectrum).
As an example, let us discuss the occurrence of negative eigenvalues for the self-adjoin extension with $\Theta=0$. It can be checked by direct inspection that the condition $\det\big{[}\Lambda(-\mu)\big{]}=0$, $\mu\in\mathbb{R}^{+}$, is fulfilled if and only if $\mu=1$ (see the explicit expression (3.9) reported in the forthcoming Lemma 3.1 for $\Lambda(z)$). Moreover, $\Lambda(-1)=0$ so that $\ker[\Lambda(-1)]=\mathbb{C}^{4}$. Accordingly, $-1\in\sigma_{\mathrm{pp}}\big{(}H_{\mathrm{P}}^{(0)}\big{)}$ is a fourthly degenerate eigenvalue of the Hamiltonian and the associated eigenspace is spanned by
$\bm{G}_{1,s}^{(\ell)}=K_{|\ell+\alpha|}(r)\,\tfrac{e^{i\ell\vartheta}}{\sqrt{2\pi}}\begin{pmatrix}\delta_{s,\uparrow}\\ \delta_{s,\downarrow}\end{pmatrix},\qquad s\in\{\uparrow,\downarrow\},\,\ell\in\{0,-1\}\,,$ (2.44)
see (3.8) below, together with (2.9) and (2.8).
@@
We do not discuss here the presence of eigenvalues embedded in the continuous spectrum. These would be exactly the exceptional points forming the set $\mathrm{e}_{+}^{(\Theta)}$ identified next in Proposition 3.3. Such a question has been investigated for regular magnetic fields in [AHK23, HH23].
@@
$f^{(\Theta,\pm)}_{(s,s^{\prime}),\mathbf{k}}:=\tfrac{e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell\in\mathbb{Z}}\left(e^{\pm i\pi|\ell+\alpha|}-e^{\pm i\pi|\ell|}\right)e^{i\ell(\vartheta-\omega_{\pm})}\,\delta_{s\,s^{\prime}}\\ +\tfrac{i\pi\,e^{\pm i\frac{\pi}{4}}}{(2\pi)^{3/2}\sqrt{k}}\sum_{\ell,\ell^{\prime}\in\{0,-1\}}\!\Big{\{}\big{[}\Lambda_{\pm}(k^{2})+\Theta\big{]}^{-1}\Big{\}}_{s^{\prime}\!,s}^{\ell^{\prime}\!,\ell}\,e^{i(\ell^{\prime}\vartheta-\ell\omega_{\pm})}\;(\pm ik)^{|\ell+\alpha|+|\ell^{\prime}+\alpha|}\,.$ (2.48)
@@
Proposition 2.21 (Zero-energy resonances for $H_{\mathrm{P}}^{(\mathrm{F})}$ and $H_{\mathrm{P}}^{(\Theta)}$).
@@
$\bm{\psi}_{0}=\begin{pmatrix}\psi_{\uparrow}\\ \psi_{\downarrow}\end{pmatrix},\qquad\psi_{s}(r,\vartheta)=\sum_{\ell\in\{0,-1\}}q_{s}^{(\ell)}\,\tfrac{2^{|\ell+\alpha|-1}\Gamma(|\ell+\alpha|)}{r^{|\ell+\alpha|}}\,\tfrac{e^{i\ell\vartheta}}{\sqrt{2\pi}},$ (2.49)